# GLOVER: Generalizable Open-Vocabulary Affordance Reasoning for Task-Oriented Grasping

### v1

@@ 4.3 Grasp Pose Estimation
We denote the gripper similarly as an ellipsoid surface following previous works [72, 42]. The gripper’s pose is defined by a 7D vector $\boldsymbol{x}=\left[x_{g},y_{g},z_{g},{q_{g}}^{x},{q_{g}}^{y},{q_{g}}^{z},{q_{g}}^{w}\right]$, where $\left(x_{g},y_{g},z_{g}\right)$ are the coordinates of the gripper’s position and $\left({q_{g}}^{x},{q_{g}}^{y},{q_{g}}^{z},{q_{g}}^{w}\right)$ are its orientation quaternions. Our pose estimation method is to identify a pose $\boldsymbol{x}$ that allows the gripper ellipsoid $\mathcal{G}$ to align with the affordance superquadric $\mathcal{A}$ while satisfying constraints that ensure $\boldsymbol{x}$ is within the gripper’s reach (Fig. 3).
@@
$\min_{\boldsymbol{x}}\sum_{i=1}^{L}\left(\left(F\left(\boldsymbol{p}_{i}^{\boldsymbol{x}},\boldsymbol{\lambda_{\mathcal{A}}}\right)-1\right)\right)^{2},$ (7)
@@
The cost function Eq.(7) aims to minimize the distance between the affordance superquadric $\mathcal{A}$ and $L$ points $\boldsymbol{p}_{i}^{\boldsymbol{x}}$, which are sampled from the closest half of the gripper ellipsoid $\mathcal{G}$. This choice prevents the gripper from penetrating the object by ensuring that only the nearest portion of $\mathcal{G}$ approaches $\mathcal{A}$, thus avoiding potential collisions. Constraint terms $C_{i}(\cdot)$ are employed to ensure the generation of safe grasping poses, such as avoiding self-collision and collision with enviromental obstacles.
Figure 4: Examples of inferred visual affordance in mulitple scenes. The testing scenes are designed to evaluate the model’s compositional object understanding (attributes, relations, complex scenes) and task-aware tools using (tool using, function reasoning).
Compositional Object Understanding Task-Aware Tool Using
Method Attribute(70) Relation(30) Complex Scene(50) Tool Use(100) Function Reason(50) #AVG
@@
VRB [3] 52.9% 37.1% 36.7% 20.0% 8.0% 6.0% 58.0% 32.0% 0.0% 0.0% 36.7% 22.3%
LERF-TOGO [65] 58.6% 45.7% 30.0% 23.3% 76.0% 62.0% 64.0% 53.0% 22.0% 10.0% 54.7% 42.7%
LISA [31] - 34.0% - 13.3% - 14.0% - 7.0% - 0.0% - 10.7%
GLOVER 95.7% 84.3% 80.0% 73.3% 82.0% 68.0% 88.0% 80.0% 76.0% 68.0% 86.0% 76.3%
Table 1: Real-world experimental results. The Attribute, Relation, Complex Scene, Tool Use and Function Reason have 7, 3, 5, 10, 5 test scenes, respectively. All the scenes are evaluated 10 times by varying the object arrangements. We report the affordance reaso ing and real-world grasping success rates to compare with previous methods.
@@ 5 Experiments
In this section, we establish a comprehensive experimental framework to validate the effectiveness and efficiency of GLOVER. We compare with previous methods in both the real-world grasping (Sec. 5.3) and visual affordance reasoning benchmark (Sec. 5.4). Efficiency analysis (Sec.5.5) and ablations (Sec. 5.6) are also performed to analyze the model components.
@@ 5.1 Implementation Details
We fine-tuned our GLOVER based on LISA-7B [31], LLaVA [39], and SAM [26] on eight NVIDIA A6000 GPUs for $5$ epochs, requiring approximately 18 hours to complete. The initial learning rate is $55e-$ 5 and we use the AdamW [43] optimizer by default. For the real-world experiments, we collect the RGB-D images with an Orbbec Femto Bolt, with an image size of $1280\times 960$. The experiments are deployed on a UR5e robot arm in our implementation.
@@ 5.2 Real-World Task Design
We evaluate GLOVER on a wide range of objects from diverse scenes, focusing on two main challenges: compositional object understanding, and task-aware tool using.
@@
We design three sub-tasks for compositional object understanding, namely attributes, relations, and complex scenes. The attributes task evaluates the agents’ understanding of colors, sizes, and materials of different objects. The relations task challenges the agents’ ability to understand spatial relations, like “above” or “below” mug shown in Fig. 4 (a). For complex scenes, we assess the model’s capacity for object-level perception by increasing the difficulty for the models to select objects based on language instructions through complex multi-object scenes (Fig. 4 (b)).
@@
For tool use, we prompt GLOVER  with phrases composed of verbs plus nouns, such as “pick up the hammer”, to assess the agent’s understanding of the general affordances of everyday tools. Examples are shown in Fig. 4 (c). In contrast, function reasoning tests the agent’s understanding of the functions associated with different parts of a tool. This task requires more precise visual perception and sophisticated reasoning, as different verbs applied to the same tool (e.g., “pick up” vs. “dispense” the sanitizer) imply varying affordances. This setup demands intricate common-sense reasoning from the agent, as shown in Fig. 4 (d).
Evaluation. The scenes and objects we used are elaborated in the Sec. 7 and Table 6. For each scene, we vary the object positions and orientations to test it 10 times, and report the success rate. The success rate includes affordance and grasping success rate, referring to the success rate of affordance reasoning and real-world grasping. For real-world grasping, we only test the scenes with successfully reasoned affordance.
@@ 5.3 Real-World Results
We report the real-world performance of GLOVER  and compare it with baseline models to demonstrate its effectiveness.
Baselines. We construct three baselines for comparisons. VRB [3] predicts contact points by learning from human video demonstrations. LERF-TOGO [65] reconstructs the scenes dynamically via LERF [25], extracting 3D object masks from DINO [4] features and conditioning object-part queries based on these masks. We also replace our GLOVER module with the original LISA-7B [31] model, which we denote as LISA, to highlight the effectiveness of our affordance fine-tuning. For a fair comparison, we provide LERF-TOGO with ambiguous language queries rather than specific part queries (which require an additional LLM to infer). Since LISA outputs binary masks, we estimate grasping poses using superquadrics based on the stereo binary mask and report the real-world grasping results.
Results. The performance are reported in Table 1. Our model surpasses the previous approach, LERF-TOGO [65], achieving an average increase of 31.3% in affordance success rate and 33.6% in real-world grasping success rate. Key findings include: (1) VRB [3] performs poorly in scenes that need reasoning based on human instructions due to the lack of language processing capabilities; (2) LERF-TOGO performs well with object recognition but lacks the complex reasoning required for interpreting object relations and diverse tool functions, likely due to the bag-of-words [62] limitation; (3) In contrast to the original LISA model, our model exhibits finer object component perception, resulting in more precise interactive regions and improved real-world grasping accuracy. In summary, GLOVER excels in both intricate common-sense reasoning and precise affordance inference.
@@ 5.4 Affordance Comparisons
Evaluation metrics. Following the previous work, we adopt the Kullback-Leibler Divergence (KLD), Similarity (SIM), and Normalized Scanpath Saliency (NSS) as evaluation metrics. Lower KLD values and higher SIM and NSS values indicate better affordance inference. Further details on these metrics are elaborated in the Sec. 8.
@@
Table 2: Visual affordance results in the benchmarking dataset. GLOVER outperforms previosu state-of-the-art methods by a large margin in all the three metrics.
@@
5.5 Efficiency Analysis
We assess GLOVER’s efficiency in terms of time required for affordance reasoning and grasping pose estimation. The process of affordance reasoning includes both scene capture and model inference. For LERF-TOGO, this process requires rendering the scene, which takes a few minutes. As shown in Table 3, GLOVER achieves approximately 330 times faster affordance reasoning compared to LERF-TOGO. For grasping pose estimation, our proposed AGE (Sec. 4) is about 40 times faster than the GraspNet [11] used in LERF-TOGO. All the time costs are reported on a single NVIDIA 4090 GPU. The results demonstrate the efficiency of our approach. This high efficiency enables GLOVER to track moving objects dynamically, deducing affordances in a real-time way. We show some real-time tracking &\&& affordance reasoning examples in Fig. 5. The qualitative results demonstrate the fast, precise, and robust tracking performance of our GLOVER.
@@ 5.6 Ablations
Camera pose. We vary the RGB-D camera poses to examine the influence of viewpoint changes. We split the camera height into Low, Mid, and High and change the horizontal view to Left, Front, and Right. Each view is tested on 30 scenes for five times. Results in Table 4 indicate minimal performance differences for horizontal view changes but notable differences for vertical view changes.
Views Low Mid High
Left 71.3% 71.3% 68.6%
Front 82.6% 85.3% 86.0%
Right 78.0% 76.6% 81.3%
Table 5: Results of GLOVER with different camera poses.
6 Conclusion & Limitations
In conclusion, our GLOVER framework demonstrates substantial advancements in open-vocabulary affordance reasoning and task-oriented grasping, outperforming existing methods in both accuracy and efficiency. By fine-tuning a foundational model with enhanced affordance understanding while preserving world knowledge, GLOVER achieves state-of-the-art performance across real-world grasping and affordance reasoning tasks. Our proposed AGE module enables rapid and precise grasping pose estimation, overcoming limitations in speed and adaptability seen in prior approaches. Extensive experiments and ablation studies confirm GLOVER’s capabilities in complex reasoning, efficient affordance inference, and robust grasping across diverse scenarios, positioning it as a versatile tool for real-world robotic manipulation.
Despite the promises, several limitations remained. First, our primary focus is on grasping tasks, with limited demonstration of effectiveness in long-horizon manipulation tasks. For long-horizon tasks, GLOVER decomposes into multiple steps, potentially resulting in less natural long-horizon behaviors. Second, the AGE cannot estimate the tactile information for the grasping, which limits the ability of grasping fragile objects like glass. In the future, we aim to improve the GLOVER from the above two aspects.
Supplementary Material
Task Sub-Tasks #Num Objects
Compositional Attribute 7 apple, mug, electric drill, pan, cup, spoon, fork
Object Relation 3 mug, cup, pan
Understanding Complex Scene 5 tape, goggles, tennis racket, electric drill, fruit
Task-Aware Tool Use 10 screwdriver, knife, scissors, hammer, pliers, saw, electric drill, tape measure, pot, pan
Tool Using Function Reason 5 knife, charger, sanitizer bottle, pot, tissue box
Table 6: The specific details of testing scenes we used in the model evaluation. Please note that our model can conduct open-vocabulary affordance reasoning and grasping on nearly any object encountered in daily life. Balancing experimental requirements and practical constraints, we selected specific objects to construct the experiments, with the aim of standardizing the testing of different models’ capabilities. (* In the Complex scene, the distractors are from all objects we own, over 5 items per scene.)
7 Real-World Evaluation Dataset
We construct 30 real-world experimental scenes, including two main tasks, namely Compositional Object Understanding and Task-Aware Tool Using.
We design three sub-tasks for compositional object understanding, namely attributes, relations, and complex scenes. The attributes task evaluates the agents’ understanding of colors, sizes, and materials of different objects. The relations task challenges the agents’ ability to understand spatial relations, like “above” or “below” mug. For complex scenes, we assess the model’s capacity for object-level perception by increasing the difficulty for the models to select objects based on language instructions through complex multi-object scenes. Note that the number of distractors in the complex scenes ranges from 5 to 15. The objects used in the task are specified in Table 6.
For tool use, we prompt GLOVER  with phrases composed of verbs plus nouns, such as “pick up the hammer”, to assess the agent’s understanding of the general affordances of everyday tools. In contrast, function reasoning tests the agent’s understanding of the functions associated with different parts of a tool. This task requires more precise visual perception and sophisticated reasoning, as different verbs applied to the same tool (e.g., “pick up” vs. “dispense” the sanitizer) imply varying affordances. This setup demands intricate common-sense reasoning from the agent. The objects we used are specified in Table 6.
8 Evaluation Metrics

### v2

@@ 4.3 Grasp Pose Estimation
We denote the gripper similarly as an ellipsoid surface following previous works [75, 45]. The gripper’s pose is defined by a 7D vector $\boldsymbol{x}=\left[x_{g},y_{g},z_{g},{q_{g}}^{x},{q_{g}}^{y},{q_{g}}^{z},{q_{g}}^{w}\right]$, where $\left(x_{g},y_{g},z_{g}\right)$ are the coordinates of the gripper’s position and $\left({q_{g}}^{x},{q_{g}}^{y},{q_{g}}^{z},{q_{g}}^{w}\right)$ are its orientation quaternions. For poses $\boldsymbol{x}$, our goal is to identify a pose $\boldsymbol{\hat{x}}$ that allows the gripper ellipsoid $\mathcal{G}$ to align with the affordance superquadric $\mathcal{A}$ while satisfying constraints that ensure $\boldsymbol{\hat{x}}$ is within the gripper’s reach (the third part in Fig. 2).
@@
$\boldsymbol{\hat{x}}=$ $\arg\min_{\boldsymbol{x}}\sum_{i=1}^{L}\left(\left(F\left(\boldsymbol{p}_{i}^{\boldsymbol{x}},\boldsymbol{\lambda_{\mathcal{A}}}\right)-1\right)\right)^{2},$ (7)
@@
The cost function Eq.(7) aims to minimize the distance between the affordance superquadric $\mathcal{A}$ and $L$ points $\boldsymbol{p}_{i}^{\boldsymbol{x}}$, while $\boldsymbol{p}_{i}^{\boldsymbol{x}}$ are sampled from the closest half of the gripper ellipsoid $\mathcal{G}$. This choice prevents the gripper from penetrating the object by ensuring that only the nearest portion of $\mathcal{G}$ approaches $\mathcal{A}$, thus avoiding potential collisions. Constraint terms $C_{i}(\cdot)$ are employed to ensure the generation of safe grasping poses, such as avoiding self-collision and collision with enviromental obstacles.
Task Sub-Tasks #Num Objects
Attribute 7 apple, mug, electric drill,
Compositional pan, cup, spoon, fork
Object Relation 3 mug, cup, pan
Understanding Complex 5 tape, goggles, tennis racket,
Scene electric drill, fruit
Tool Use 10 screwdriver, knife, scissors,
Task-Aware hammer, pliers, electric drill,
Tool Using tape measure, saw, pot, pan
Function 5 knife, charger, sanitizer
Reason bottle, pot, tissue box
Table 1: The specific details of testing scenes we used in the model evaluation. GLOVER can conduct open-vocabulary affordance reasoning and grasping on nearly any object encountered in daily life. Balancing experimental requirements and practical constraints, we selected specific objects to construct the experiments, with the aim of standardizing the testing of different models’ capabilities. (* In the Complex Scene, the distractors are from all objects we own, over 5 items per scene.)
Figure 3: Eval Examples: Examples of inferred visual affordance and grasping pose in multiple scenes. The testing scenes are designed to evaluate the model’s compositional object understanding (attributes, relations, complex scenes) and task-aware tools using (tool using, function reasoning). Generalization: GLOVER presents open-vocabulary ability across diverse environments, including real-world, simulator (RLBench [24]), scenes from other datasets (Ego4D [17]), and across-embodiments (humanoid robots with dexterous hands).
Compositional Object Understanding (%) Task-Aware Tool Using (%)
Method Attribute(70) Relation(30) Complex Scene(50) Tool Use(100) Function Reason(50) #AVG (%)
@@
VRB [3] 52.9 37.1 36.7 20.0 8.0 6.0 58.0 32.0 0.0 0.0 36.7 22.3
LERF-TOGO [68] 58.6 45.7 30.0 23.3 76.0 62.0 64.0 53.0 22.0 10.0 54.7 42.7
LISA [33] - 34.0 - 13.3 - 14.0 - 7.0 - 0.0 - 10.7
RAM [32] 64.3 57.1 43.3 26.7 78.0 62.0 84.0 62.0 34.0 12.0 66.0 49.0
GLOVER 95.7 84.3 80.0 73.3 82.0 68.0 88.0 80.0 76.0 68.0 86.0 76.3
Table 2: Real-world experimental results. The Attribute, Relation, Complex Scene, Tool Use and Function Reason have 7, 3, 5, 10, 5 test scenes, respectively. All the scenes are evaluated 10 times by varying the object arrangements. We report the affordance reasoning and real-world grasping success rates to compare with previous methods.
@@ 5 Experiments
In this section, we establish a comprehensive experimental framework to validate the effectiveness and efficiency of GLOVER. We compare with previous methods in both the real-world grasping (Sec. 5.3) and visual affordance reasoning benchmark (Sec. 5.4). Ablations (Sec. 4) are also performed to analyze the model components.
@@ 5.1 Implementation Details
We fine-tuned our GLOVER based on LISA-7B [33], LLaVA [42], and SAM [28] on eight NVIDIA A6000 GPUs for $5$ epochs, requiring approximately 18 hours to complete. The initial learning rate is $55e-$ 5 and we use the AdamW [46] optimizer by default. For the real-world experiments, we collect the RGB-D images with an Orbbec Femto Bolt, with an image size of $1280\times 960$ for UR5e robot arm. The humanoid robot, Unitree G1, is utilized for cross-embodiment generalization with more real downstream experiments. The G1 robot is equipped with Inspire dexterous hands RH56DFX, and we use the original Intel RealSense D435i of the robot to capture RGB-D images with the size of $640\times 480$. The details can be found in supplementary material.
@@ 5.2 Real-World Task Design
We evaluate GLOVER on a wide range of objects from diverse scenes, focusing on two main challenges: compositional object understanding, and task-aware tool using. The scenes and objects we used are elaborated in the Table 1.
@@
We design three sub-tasks for compositional object understanding, namely attributes, relations, and complex scenes. The attributes task evaluates the agents’ understanding of colors, sizes, and materials of different objects. The relations task challenges the agents’ ability to understand spatial relations, like “above” or “below” mug shown in Fig. 3. For complex scenes, we assess the model’s capacity for object-level perception by increasing the difficulty for the models to select objects based on language instructions through complex multi-object scenes.
@@
For tool use, we prompt GLOVER  with phrases composed of verbs plus nouns, such as “pick up the hammer”, to assess the agent’s understanding of the general affordances of everyday tools. In contrast, function reasoning tests the agent’s understanding of the functions associated with different parts of a tool. This task requires more precise visual perception and sophisticated reasoning, as different verbs applied to the same tool (e.g., “pick up” vs. “pump” the sanitizer) imply varying affordances. This setup demands intricate common-sense reasoning from the agent, as shown in Fig. 3.
Evaluation. For each scene, we vary the object positions and orientations to test it 10 times, and report the success rate. The success rate includes affordance and grasping success rate, referring to the success rate of affordance reasoning and real-world grasping. We manually defined the ground truth regions for different objects based on the language instructions of the required task. As long as the majority of filtered affordance points fall within these ground truth regions, it is considered an affordance success. Grasping success is defined as the ability to complete real-world grasping. For real-world grasping, we only test the scenes with successfully reasoned affordance.
@@ 5.3 Real-World Results
Baselines. We construct four baselines for comparisons. VRB [3] predicts contact points by learning from human video demonstrations. LERF-TOGO [68] reconstructs the scenes dynamically via LERF [27], extracting 3D object masks from DINO [4] features and conditioning object-part queries based on these masks. We also replace our GLOVER module with the original LISA-7B [33] model, which we denote as LISA, to highlight the effectiveness of our affordance fine-tuning. RAM [32] constructs the affordance memory from 2D images and retrieves similar demonstrations from the affordance memory with the help of CLIP [66] to reason affordance in the unseen domain. For a fair comparison, we provide LERF-TOGO with ambiguous language queries rather than specific part queries (which require an additional LLM to infer). Since LISA outputs binary masks, we estimate grasping poses using superquadrics based on the stereo binary mask and report the real-world grasping results.
Results. The performance are reported in Table 2. Our model surpasses the previous approach, RAM [32], achieving an average increase of 20.0% in affordance success rate and 17.3% in real-world grasping success rate. Key findings include: (1) VRB [3] performs poorly in scenes that need reasoning based on human instructions due to the lack of language processing capabilities; (2) LERF-TOGO performs well with object recognition but lacks the complex reasoning required for interpreting object relations and diverse tool functions, likely due to the bag-of-words [65] limitation; (3) In contrast to the original LISA model, our model exhibits finer object component perception, resulting in more precise interactive regions and improved real-world grasping accuracy. In summary, GLOVER excels in both intricate common-sense reasoning and precise affordance inference.
@@ 5.4 Affordance Comparisons
Evaluation metrics. Following the previous work, we adopt the Kullback-Leibler Divergence (KLD), Similarity (SIM), and Normalized Scanpath Saliency (NSS) as evaluation metrics. Lower KLD values and higher SIM and NSS values indicate better affordance inference. Details on these metrics are elaborated in the supplementary material.
@@
Table 3: Visual affordance results in the benchmarking dataset. GLOVER outperforms previous state-of-the-art methods by a large margin in all three metrics.
@@
5.5 Ablations
Efficiency analysis. We assess GLOVER’s efficiency in terms of time required for affordance reasoning and grasping pose estimation. The process of affordance reasoning includes both scene capture and model inference. As shown in Table 4, GLOVER achieves approximately 330 times and 29 times faster affordance reasoning compared to LERF-TOGO [68] and RAM [32]. For grasping pose estimation, our proposed AGE (Sec. 4) is about 40 times faster than the GraspNet [11] used in LERF-TOGO and RAM. All the time costs are reported on a single NVIDIA 4090 GPU. The results demonstrate the efficiency of our approach. This high efficiency enables GLOVER to track moving objects dynamically, deducing affordances in a real-time way. The tracking performance can be found in the supplementary materials.
@@ 5.5 Ablations
Generalization. We show the visualization of open-vocabulary affordance reasoning in diverse scenarios to show the generalization in Fig. 3, including real-world, simulator, other video datasets. We also validate GLOVER’s generalization ability across embodiments, showing the effectiveness of the proposed method in humanoid robots with dexterous hands. The Inverse Kinematics is utilized to determine the motions for reaching the target grasping pose output by GLOVER. We conduct experiments on four tasks, open oven, pump sanitizer, pick up sanitizer, and pick up mug. The results are shown in Table 6. More details and failure cases are elaborated in the supplementary material.
Task Open Pump Pick up Pick up
Oven Sanitizer Sanitizer Mug
GLOVER 4/5 3/5 2/5 3/5
Table 6: Success rate of GLOVER in humanoid robot with dexterous hands.
6 Conclusion
In conclusion, our GLOVER framework demonstrates substantial advancements in open-vocabulary affordance reasoning and task-oriented grasping, outperforming existing methods in both accuracy and efficiency. By fine-tuning a foundational model with enhanced affordance understanding while preserving world knowledge, GLOVER achieves state-of-the-art performance across real-world grasping and affordance reasoning tasks. The AGE module enables rapid and precise grasping pose estimation in a non-parametric manner. Extensive experiments and ablation studies confirm GLOVER’s capabilities in complex reasoning, efficient affordance inference, and robust grasping across diverse scenarios and embodiments.
Appendix A Evaluation Metrics
@@
Appendix B Real-World Experiment Settings
We introduce the settings of our real-world experiments, including the single table-top robotic arm experiments and the humanoid robot experiments.
For the single table-top robotic arm, we utilize a UR5e robotic arm equipped with a DH PGI gripper and set an Orbbec Femto Bolt on the front side of the workspace. The default image size of the captured RGB-D stream is $1280\times 960$. The setting is shown as Fig. 4.
For the humanoid robot, we use the Unitree G1 equipped with Inspire RH56DFX dexterous hands and Intel RealSense D435i RGB-D camera. The image size for the RGB-D images is set as $640\times 480$. The setting is shown as Fig. 5.
More experiments can be found in Fig. 7 and the video in supplementary material.
Figure 4: The experiment setting of single table-top robotic arm.
Figure 5: The experiment setting of a humanoid robot with dexterous hands.
Appendix C Failure Cases
The failure cases in the real-robot experiments can be summarized as two main points:
  • We use the Inverse Kinematics (IK) to calculate the motions based on the target pose, which leads to rough grasping motions. The lack of dexterous grasping policy leads to collisions and overturns, as shown in the first and second row of Fig. 6. In the future, we aim to train ACT or Diffusion Policy for the grasping motion optimization.
  • For the calibration of the relationship between the hand and eye of humanoid robots, our measurements may have certain errors, which could lead to the failure of tasks requiring fine operations, as shown in the third row of Fig. 7.
Figure 6: Failure cases. Row 1 & 2: Rough grasping motions lead to collisions and overturns. Row 3: The cumulative error in hand-eye calibration leads to inaccurate positions.
Appendix D Tracking Performance
The real-time performance of our GLOVER makes tracking moving objects while reasoning affordance possible. Time cost can be found in Table 3 in the main manuscript. We show some real-time tracking &\&& affordance reasoning examples in Fig. 8. The qualitative results demonstrate the fast, precise, and robust tracking performance of our GLOVER. More examples can be found in the video of supplementary material.
Figure 7: Demonstrations of manipulation tasks completed by the Unitree G1.